Mr. Azu invested an amount at rate of 12% per annum and invested another amount, GH¢
580.00 more than the first at 14%. If Mr. Azu had total accumulated amount of
GH¢2,358.60, how much was his total investment?
1
Expert's answer
2020-06-11T17:52:45-0400

Let total amount invested be x at rate 12%.

so amount invested at rate 14% will be x+580.00


total account collected at end of year is 2358.6

2358.6 GHc


So

"x(1.12) + (x+580.0)(1.14) = 2358.6"

solving this,

we get 

"1.12x+1.14x +580*1.14 = 2358.6"

"2.26x + 661.2=2358.6"

"2.26x = 2358.6 - 661.2"

"2.26x = 1697.4"


"x=751.06"  GHc